OpenStudy (anonymous):

in parallelogram HOPE. HO=3x-2 EP= 2x and PO=x+12. find the perimeter

OpenStudy (anonymous):

H0=EP 3x-2=2x\\ x=2 Can you finish it?

OpenStudy (anonymous):

HO= 4 EP=4 PO=14 HE+14 perimeter= 14 + 14 + 4 + 4= 36

OpenStudy (anonymous):

thank you :)